A two-digit number is 1 more than eight times the sum of the digits. The ones digit is 3 less than the ten digit. What is the number?
Answer by swincher4391(1107)

10x + y = 8(x+y) + 1
y = x-3
SYSTEM OF EQUATIONS
10x + y = 8x + 8y +1
2x - 7y = 1
-x + y = -3
2x - 7y = 1
-2x + 2y = -6
---------------
-5y = -5
y = 1
1 = x-3
x = 4
The number is 41.